OpenStudy (anonymous):

Given 3x = 4y = 12z, show that z = xy/(x + y) .

OpenStudy (anonymous):

can u give me a minute?

OpenStudy (anonymous):

oops \[3^x=3^y=12^z z=\frac{xy}{x+a} \]

OpenStudy (anonymous):

x+y

OpenStudy (anonymous):

let: \[3^x = 4^y = 12^z = k\] \[3 = k^{1/x}\] \[4 = k^{1/y}\] \[12 = k^{1/z}\] Use: \[3 \times 4 = 12\] \[k^{1/x} \times k^{1/y} = k^{1/z}\] \[\implies \frac{1}{x} + \frac{1}{y} = \frac{1}{z}\] \[\frac{1}{z} = \frac{x + y}{xy}\] \[\implies \color{blue}{z = \frac{xy}{x+y}}\]

OpenStudy (anonymous):

wow @waterineyes u r faster than me

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

when i solved it u already posted the solution

OpenStudy (anonymous):

ohh can u solve using log?

OpenStudy (anonymous):

I can try..

OpenStudy (anonymous):